# Can Watermarked LLMs be Identified by Users via Crafted Prompts?

### v1

@@ Appendix A Proof of Theorem 1 and Additional Analyses
Goal: Prove that under the Lipschitz continuity condition, the watermark modification differences for two similar prompts $p_{1}$ and $p_{2}$ have high similarity, with an expected similarity of at least $\rho$.
@@
For a prompt $p$ and keys $k_{1},k_{2}$, define the watermark modification difference as:
$\Delta_{p}(k_{1},k_{2})=P_{M}^{F}(\cdot|p,k_{1})-P_{M}^{F}(\cdot|p,k_{2})$ (13)
@@
$\|P_{M}^{F}(\cdot|p_{1},k)-P_{M}^{F}(\cdot|p_{2},k)\|_{1}\leq L\cdot\|P_{M}(\cdot|p_{1})-P_{M}(\cdot|p_{2})\|_{1}$ (14)
Since $p_{1}$ and $p_{2}$ satisfy the similarity condition $\text{KL}(P_{M}(\cdot|p_{1})||P_{M}(\cdot|p_{2}))\leq\epsilon$ (Equation 5), by Pinsker’s inequality (Csiszár & Körner, 2011):
$\|P_{M}(\cdot|p_{1})-P_{M}(\cdot|p_{2})\|_{1}\leq\sqrt{2\epsilon}$ (15)
@@
$\|P_{M}^{F}(\cdot|p_{1},k)-P_{M}^{F}(\cdot|p_{2},k)\|_{1}\leq L\cdot\sqrt{2\epsilon}$ (16)
@@
Consider the definition of $\Delta_{p}(k_{1},k_{2})$ and compute $\|\Delta_{p_{1}}(k_{1},k_{2})-\Delta_{p_{2}}(k_{1},k_{2})\|_{1}$:
$\|\Delta_{p_{1}}(k_{1},k_{2})-\Delta_{p_{2}}(k_{1},k_{2})\|_{1}$ $=\|(P_{M}^{F}(\cdot|p_{1},k_{1})-P_{M}^{F}(\cdot|p_{1},k_{2}))-(P_{M}^{F}(\cdot|p_{2},k_{1})-P_{M}^{F}(\cdot|p_{2},k_{2}))\|_{1}$
$=\|(P_{M}^{F}(\cdot|p_{1},k_{1})-P_{M}^{F}(\cdot|p_{2},k_{1}))-(P_{M}^{F}(\cdot|p_{1},k_{2})-P_{M}^{F}(\cdot|p_{2},k_{2}))\|_{1}$
$\leq\|P_{M}^{F}(\cdot|p_{1},k_{1})-P_{M}^{F}(\cdot|p_{2},k_{1})\|_{1}+\|P_{M}^{F}(\cdot|p_{1},k_{2})-P_{M}^{F}(\cdot|p_{2},k_{2})\|_{1}$
@@
Assume $\text{Sim}(\cdot,\cdot)$ is cosine similarity, which is negatively correlated with $L_{1}$ distance. Since $\|\Delta_{p_{1}}(k_{1},k_{2})-\Delta_{p_{2}}(k_{1},k_{2})\|_{1}\leq\delta^{\prime}$, and $\delta^{\prime}$ is a small positive number:
$\text{Sim}(\Delta_{p_{1}}(k_{1},k_{2}),\Delta_{p_{2}}(k_{1},k_{2}))\geq\rho^{\prime}$ (18)
@@
Since $\text{Sim}(\Delta_{p_{1}}(k_{1},k_{2}),\Delta_{p_{2}}(k_{1},k_{2}))\geq\rho^{\prime}$ for any $k_{1},k_{2}\in\mathcal{K}$, for randomly sampled $k_{1},k_{2}$:
$\mathbb{E}_{k_{1},k_{2}\sim\mathcal{K}}\left[\text{Sim}(\Delta_{p_{1}}(k_{1},k_{2}),\Delta_{p_{2}}(k_{1},k_{2}))\right]\geq\rho^{\prime}$ (19)
@@
Conclusion: Under the Lipschitz continuity condition, for two similar prompts $p_{1}$ and $p_{2}$, the expected similarity of their watermark modification differences under randomly sampled watermark keys $k_{1}$ and $k_{2}$ is at least $\rho$, where $\rho$ is a large positive number close to 1. This proves the high consistency of watermark effects in similar contexts, ensuring the detectability and robustness of the watermark. ∎
@@ Appendix B Statistical Analysis of Unwatermarked LLMs Identification
For an unwatermarked model, $P_{M}^{F}=P_{M}$ for any $k$. Therefore:
$\Delta_{R}(P_{i},k_{m},k_{n})$ $=R(\hat{P}_{M}^{F}(\cdot|P_{i},k_{m}))-R(\hat{P}_{M}^{F}(\cdot|P_{i},k_{n}))$ (20)
$=R(\hat{P}_{M}(\cdot|P_{i}))-R(\hat{P}_{M}(\cdot|P_{i}))$
@@
Consequently, $\text{Sim}(\Delta_{R}(P_{i},k_{m},k_{n}),\Delta_{R}(P_{j},k_{m},k_{n}))\approx\text{Sim}(0,0)=0$ for any $P_{i},P_{j},k_{m},k_{n}$. This results in the average similarity $\bar{S}$ in Equation 9 being approximately 0.
@@ Appendix E Detection of Global-Fixed Key Watermarking in LLMs
Watermark-Probe-1 (w. prompt 4.3)
@@
Watermark-Probe-2 (w. prompt 4.3)
@@
Watermark-Probe-v1 (w. prompt 4.3)
@@
Watermark-Probe-v2 (w. prompt 4.3)
@@ Appendix F Details of Tested Watermarking Algorithms
To help understand the watermarking algorithms related to the experiments in this paper, we provide detailed information for all watermarking algorithms in Table 5, including their names, references, types, and brief descriptions. All our experiments were conducted using the MarkLLM (Pan et al., 2024a) framework.

### v3

@@ 7 Conclusion
8 Acknowledgements
This work is primarily supported by the Key Research and Development Program of China (No. 2024YFB3309702). Additional support was provided by the National Science Foundation (NSF) under grants III-2106758 and POSE-2346158.
We would like to express our gratitude to the anonymous ICLR 2025 reviewers (Reviewer GLo8, DykZ, qXQr, AeYS) and Area Chair R3na for their valuable feedback and suggestions that helped improve this paper. We also thank Nikola Jovanović for his insightful public comments.
@@ Appendix A Proof of Theorem 1 and Additional Analyses
Goal: Prove that under the Lipschitz continuity condition, the watermark modification differences for two similar prompts $x_{1}$ and $x_{2}$ have high similarity, with an expected similarity of at least $\rho$.
@@
For a prompt $x$ and keys $k_{1},k_{2}$, define the watermark modification difference as:
$\Delta_{x}(k_{1},k_{2})=P_{M}^{F}(\cdot|x,k_{1})-P_{M}^{F}(\cdot|x,k_{2})$ (13)
@@
$\|P_{M}^{F}(\cdot|x_{1},k)-P_{M}^{F}(\cdot|x_{2},k)\|_{1}\leq L\cdot\|P_{M}(\cdot|x_{1})-P_{M}(\cdot|x_{2})\|_{1}$ (14)
Since $x_{1}$ and $x_{2}$ satisfy the similarity condition $\text{KL}(P_{M}(\cdot|x_{1})||P_{M}(\cdot|x_{2}))\leq\epsilon$ (Equation 5), by Pinsker’s inequality (Csiszár & Körner, 2011):
$\|P_{M}(\cdot|x_{1})-P_{M}(\cdot|x_{2})\|_{1}\leq\sqrt{2\epsilon}$ (15)
@@
$\|P_{M}^{F}(\cdot|x_{1},k)-P_{M}^{F}(\cdot|x_{2},k)\|_{1}\leq L\cdot\sqrt{2\epsilon}$ (16)
@@
Consider the definition of $\Delta_{x}(k_{1},k_{2})$ and compute $\|\Delta_{x_{1}}(k_{1},k_{2})-\Delta_{x_{2}}(k_{1},k_{2})\|_{1}$:
$\|\Delta_{x_{1}}(k_{1},k_{2})-\Delta_{x_{2}}(k_{1},k_{2})\|_{1}$ $=\|(P_{M}^{F}(\cdot|x_{1},k_{1})-P_{M}^{F}(\cdot|x_{1},k_{2}))-(P_{M}^{F}(\cdot|x_{2},k_{1})-P_{M}^{F}(\cdot|x_{2},k_{2}))\|_{1}$
$=\|(P_{M}^{F}(\cdot|x_{1},k_{1})-P_{M}^{F}(\cdot|x_{2},k_{1}))-(P_{M}^{F}(\cdot|x_{1},k_{2})-P_{M}^{F}(\cdot|x_{2},k_{2}))\|_{1}$
$\leq\|P_{M}^{F}(\cdot|x_{1},k_{1})-P_{M}^{F}(\cdot|x_{2},k_{1})\|_{1}+\|P_{M}^{F}(\cdot|x_{1},k_{2})-P_{M}^{F}(\cdot|x_{2},k_{2})\|_{1}$
@@
Assume $\text{Sim}(\cdot,\cdot)$ is cosine similarity, which is negatively correlated with $L_{1}$ distance. Since $\|\Delta_{x_{1}}(k_{1},k_{2})-\Delta_{x_{2}}(k_{1},k_{2})\|_{1}\leq\delta^{\prime}$, and $\delta^{\prime}$ is a small positive number:
$\text{Sim}(\Delta_{x_{1}}(k_{1},k_{2}),\Delta_{x_{2}}(k_{1},k_{2}))\geq\rho^{\prime}$ (18)
@@
Since $\text{Sim}(\Delta_{x_{1}}(k_{1},k_{2}),\Delta_{x_{2}}(k_{1},k_{2}))\geq\rho^{\prime}$ for any $k_{1},k_{2}\in\mathcal{K}$, for randomly sampled $k_{1},k_{2}$:
$\mathbb{E}_{k_{1},k_{2}\sim\mathcal{K}}\left[\text{Sim}(\Delta_{x_{1}}(k_{1},k_{2}),\Delta_{x_{2}}(k_{1},k_{2}))\right]\geq\rho^{\prime}$ (19)
@@
Conclusion: Under the Lipschitz continuity condition, for two similar prompts $x_{1}$ and $x_{2}$, the expected similarity of their watermark modification differences under randomly sampled watermark keys $k_{1}$ and $k_{2}$ is at least $\rho$, where $\rho$ is a large positive number close to 1. This proves the high consistency of watermark effects in similar contexts, ensuring the detectability and robustness of the watermark. ∎
@@ Appendix B Statistical Analysis of Unwatermarked LLMs Identification
For an unwatermarked model, $P_{M}^{F}=P_{M}$ for any $k$. Let $\hat{R}_{N}^{(1)}(\cdot)$ and $\hat{R}_{N}^{(2)}(\cdot)$ denote two independent empirical estimates of $R(\cdot)$ using $N$ samples each (because in the black-box setting, we cannot directly access the true $P_{M}^{F}$). Therefore:
$\Delta_{R}(x_{i},k_{m},k_{n})$ $=\hat{R}_{N}^{(1)}(P_{M}^{F}(\cdot|x_{i},k_{m}))-\hat{R}_{N}^{(2)}(P_{M}^{F}(\cdot|x_{i},k_{n}))$ (20)
$=\hat{R}_{N}^{(1)}(P_{M}(\cdot|x_{i}))-\hat{R}_{N}^{(2)}(P_{M}(\cdot|x_{i}))$
$=[\hat{R}_{N}^{(1)}(P_{M}(\cdot|x_{i}))-R(P_{M}(\cdot|x_{i}))]-[\hat{R}_{N}^{(2)}(P_{M}(\cdot|x_{i}))-R(P_{M}(\cdot|x_{i}))]$
$=\epsilon_{1}-\epsilon_{2}$
where $\epsilon_{1},\epsilon_{2}$ represent independent sampling errors that follow normal distributions $\mathcal{N}(0,\sigma^{2})$ according to the Central Limit Theorem.
Consequently, for any $x_{i},x_{j},k_{m},k_{n}$:
$\text{Sim}(\Delta_{R}(x_{i},k_{m},k_{n}),\Delta_{R}(x_{j},k_{m},k_{n}))=\text{Sim}(\epsilon_{1}-\epsilon_{2},\epsilon_{3}-\epsilon_{4})$ (21)
where $\epsilon_{1},\epsilon_{2}$ represent independent sampling errors that follow normal distributions $\mathcal{N}(0,\sigma^{2})$ according to the Central Limit Theorem.
Consequently, for any $x_{i},x_{j},k_{m},k_{n}$:
$\text{Sim}(\Delta_{R}(x_{i},k_{m},k_{n}),\Delta_{R}(x_{j},k_{m},k_{n}))=\text{Sim}(\epsilon_{1}-\epsilon_{2},\epsilon_{3}-\epsilon_{4})$ (22)
where $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}$ are independent sampling errors. To understand why this similarity has an expected value of zero, recall that cosine similarity is defined as:
$\text{Sim}(a,b)=\frac{a\cdot b}{||a||\cdot||b||}$ (23)
Note that $(\epsilon_{1}-\epsilon_{2})$ and $(\epsilon_{3}-\epsilon_{4})$ are differences of independent normal variables, each following $\mathcal{N}(0,2\sigma^{2})$. For the numerator:
$E[(\epsilon_{1}-\epsilon_{2})(\epsilon_{3}-\epsilon_{4})]$ $=E[\epsilon_{1}\epsilon_{3}]-E[\epsilon_{1}\epsilon_{4}]-E[\epsilon_{2}\epsilon_{3}]+E[\epsilon_{2}\epsilon_{4}]$ (24)
$=E[\epsilon_{1}]E[\epsilon_{3}]-E[\epsilon_{1}]E[\epsilon_{4}]-E[\epsilon_{2}]E[\epsilon_{3}]+E[\epsilon_{2}]E[\epsilon_{4}]$
@@
where the second equality follows from independence, and the final equality holds because $E[\epsilon_{i}]=0$ for all $i$. As these differences are independent and orthogonal in expectation, their cosine similarity has an expected value of zero:
$\mathbb{E}[\text{Sim}(\epsilon_{1}-\epsilon_{2},\epsilon_{3}-\epsilon_{4})]=0$ (25)
Therefore, the average similarity $\bar{S}$ in Equation 9 has an expected value of 0.
@@ Appendix E Detection of Global-Fixed Key Watermarking in LLMs
Watermark-Probe-1 (w. prompt 3.3)
@@
Watermark-Probe-2 (w. prompt 3.3)
@@
Watermark-Probe-v1 (w. prompt 3.3)
@@
Watermark-Probe-v2 (w. prompt 3.3)
@@ Appendix F Details of Tested Watermarking Algorithms
To help understand the watermarking algorithms related to the experiments in this paper, we provide detailed information for all watermarking algorithms in Table 5, including their names, references, types, and brief descriptions. All our experiments were conducted using the MarkLLM (Pan et al., 2024) framework.
@@ Appendix H Supplementary Z-Scores and P-values
Appendix I Comparison with Related Work (Gloaguen et al., 2024)
  1. 1.
    Universal Detection: Our method (particularly Water-Probe-v2) can detect all current watermarking-during-generation approaches (those that modify generation logits or sampling processes). In contrast, Gloaguen et al. (2024)’s method requires specific designs for different watermarking algorithms:
    • Monte Carlo permutation test for red-green watermarking
    • Mann-Whitney U test for EXP-edit watermarking
    • Potential new methods for future watermarking methods
    Our approach represents the first universal detection method effective across all current LLM watermarking techniques.
  2. 2.
    Unified Theoretical Foundation: We provide a unified theoretical analysis and explanation for why watermarked LLMs can be detected, specifically demonstrating how watermark key conflicts lead to identifiable characteristics in model outputs. This theoretical framework provides a comprehensive understanding of the detection mechanism.
  3. 3.
    Imperceptibility Enhancement: Beyond detection methods, we also contribute the Water-Bag approach for improving the imperceptibility of watermarked LLMs, demonstrating significant improvements in watermark concealment while maintaining detectability.
  4. 4.
    Broader Applicability for Challenging Watermarking Variants: Our method supports more challenging watermarking variants. For instance, while Gloaguen et al. (2024)’s experiments with EXP-edit only considered argmax sampling after exponential transformation (limiting a length-N watermark key list to at most N different sampling results), our method requires no such assumptions.
To demonstrate the broader applicability of our method, we conducted experiments with EXP-edit using sampling after exponential transformation. Given logits $l_{i}$, we first compute probabilities $p_{i}$ through temperature scaling:
$p_{i}=\frac{\exp(l_{i}/\tau)}{\sum_{j}\exp(l_{j}/\tau)}$ (28)
where $\tau$ is the temperature parameter. While Gloaguen et al. (2024)’s analysis focused on the deterministic argmax sampling variant:
$i^{*}=\arg\max_{i}(\xi_{i}^{(j)})^{1/p_{i}}$ (29)
This deterministic approach has a fundamental limitation - for a watermark key list of length N, it can only produce at most N distinct outputs. We instead evaluate multinomial sampling from the distribution:
$P(i)\propto(\xi_{i}^{(j)})^{1/p_{i}}$ (30)
Table 9: Experimental Results with Argmax Sampling On Exp-Edit
Temp Watermark F1 Perplexity Gloaguen P-value Water-Probe-v2 P-value
0.2 0.664 11.8 <0.001 <0.001
0.3 0.666 11.5 <0.001 <0.001
0.4 0.678 11.2 <0.001 <0.001
0.5 0.793 10.9 <0.001 <0.001
0.6 0.907 10.7 <0.001 <0.001
0.7 0.965 10.5 <0.001 <0.001
0.8 0.987 10.4 <0.001 <0.001
0.9 0.987 10.3 <0.001 <0.001
1.0 0.987 10.3 <0.001 <0.001
Table 10: Experimental Results with Multinomial Sampling On Exp-Edit, which is the most challenging case for watermarked LLM identification.
Temp Watermark F1 Perplexity Gloaguen P-value Water-Probe-v2 P-value
0.2 0.666 11.1 <0.001 <0.001
0.3 0.662 11.8 0.33 <0.001
0.4 0.672 11.5 0.83 <0.001
0.5 0.740 11.2 1.0 <0.001
0.6 0.877 11.0 1.0 <0.001
0.7 0.985 10.8 1.0 <0.001
0.8 0.985 10.7 1.0 <0.001
0.9 0.987 10.6 1.0 <0.001
1.0 0.987 10.6 1.0 <0.001
Our experiments used the MarkLLM framework with EXP-Edit watermarking (key length = 420), using OPT-1.3B as the base model and LLaMA-7B for perplexity calculations. For Gloaguen et al. (2024)’s testing method, we generated 1,000 text samples of length 200 tokens each. For Water-Probe-v2 testing, we generated 10,000 text samples of length 5 tokens each.
Tables 9 and 10 show that when applying multinomial sampling after EXP transformation, the watermark detection F1 scores and perplexity values remain largely unaffected. However, while our method maintains its effectiveness, Gloaguen et al. (2024)’s method fails to detect the watermark. This demonstrates the broader applicability of our approach to practical watermarking deployments.
Appendix J Guidelines for Constructing Watermarked LLM Identification Prompts
To help better understand our method, we provide guidelines for constructing watermarked LLM identification prompts, divided into three parts: question space design, answer space design, and implementation and verification protocols.
J.1 Overall Prompt Structure
The prompts in our identification method consist of two essential components - a prefix component and a question component. The specific requirements for each component will be detailed in subsequent sections. Here is a basic illustration of the structure:
Basic Two-Component Prompt Structure Input Prompt: Please start your answer with "WXYZ" (prefix component) and then answer the question: What is a major city in Asia? (question component) Response: WXYZ Tokyo Explanation: The generated prefix would help to fix the watermark key, while the actual answer would reflect the model’s response distribution (achieved by repeated sampling).
J.2 Question Component Design
As described in Section 3.2 Step 1, we should first construct highly correlated prompts with significantly overlapping but non-identical answer spaces. This enables easy assessment of how potential watermark keys affect the answer spaces of different prompts.
Here is a list of criteria for selecting questions:
  1. 1.
    Answer Space Similarity: Select questions with overlapping but non-identical answer spaces. For example:
    • “Name a country with a large population”
    • “Name a country with a large area”
    • “Name a country with a high GDP”
    • “Name a country with rich natural resources”
    These questions typically share common answers (e.g., USA, China, Russia) while maintaining distinct probability distributions over the answer space.
  2. 2.
    Structural Requirements:
    • Questions should be concise and unambiguous
    • Answers should come from a well-defined finite set (e.g., countries, cities)
    • Questions should maintain comparable difficulty levels
    • The target entity category should remain consistent within a test suite
J.3 Prefix Component Design
J.3.1 Watermark-Probe-v1 Construction
For Watermark-Probe-v1, simply instruct the LLM to generate a fixed prefix before answering the question through explicit prompt instructions.
Here are the design principles for the prefix component:
  1. 1.
    Use meaningless character sequences (e.g., “abcd”, “wxyz”) that have no semantic meaning in any language
  2. 2.
    Avoid any sequences that could form acronyms, abbreviations or meaningful patterns
  3. 3.
    Ensure the prefix is completely unrelated to any potential answers or question domains
  4. 4.
    Keep the prefix length sufficient for determining the watermark key while maintaining semantic independence
Here is an example of the prefix component:
  • Implementation Example for Watermark-Probe-v1 Please generate abcd before answering the question. Question: Name a country with a large population. Answer: abcd India Explanation: The generated prefix is meaningless and unrelated to the question domain, ensuring that it does not introduce any contextual bias.
J.3.2 Watermark-Probe-v2 Construction
For Watermark-Probe-v2, we need to design a controlled randomization process before answering the question to help fix the watermark key (see Section 3.3 for detailed reasons).
Here are the design principles for the prefix component:
  1. 1.
    Ensure the prefix generation does not influence the answer to the main question
  2. 2.
    Design multiple choice sets with logically equivalent probabilities
  3. 3.
    Keep the number of choices moderate and manageable
  4. 4.
    Maintain clear boundaries between different choice sets
Here is an example of the prefix component:
  • Implementation Example for Watermark-Probe-v2 Please generate a sentence that satisfies the following conditions: First word: Randomly sampled from A-Z Second word: Randomly sampled from zero to nine Third word: Randomly sampled from {cat, dog, tiger, lion} Then answer: Name a country with a large population. Answer: A one cat China Explanation: All the possible generated prefixes are not related to the question domain, ensuring that they do not introduce any contextual bias.
Appendix K Threat Model
In this section, we outline the threat model under which our watermark identification method (detector) operates. We consider the capabilities and limitations of both the detector and the LLM service provider.
K.1 Detector Capabilities
We assume the detector:
  • Has black-box access to the LLM through standard API interfaces
  • Can only interact with the model through normal prompt-response queries
  • Has no access to model architecture, parameters, or training data
  • Can perform multiple queries
  • Cannot modify or influence the model’s internal state
K.2 Trust Assumptions
The threat model assumes:
  • The LLM service provider may embed watermarks in the model outputs
  • The API interface itself is trustworthy and returns genuine model outputs
  • No man-in-the-middle attacks or response tampering occurs
  • The detection process does not require knowledge of specific watermarking algorithms
K.3 Detection Goals and Constraints
The primary objectives within this threat model are:
  • Determine the presence or absence of watermarks in model outputs
  • Maintain detection accuracy across different sampling temperatures and model configurations
Key constraints include:
  • Detection must be performed solely through black-box testing
  • Watermark removal or tampering is outside the scope
  • Detection methods must be robust against normal model output variations
Appendix L Test on Closed-source Models
We evaluated Water-Probe-V2’s detection capabilities on several closed-source models, including GPT-4o-mini, GPT-4o, GPT-3.5-turbo, Gemini-1.5-flash, and Gemini-1.5-pro. For all experiments, we utilized the latest API versions of these models (as of November 15, 2024) with a temperature setting of 0.7.
Table 11: Watermarked LLM Identification Results on Closed-source Models
Model Similarity Std Dev Z-score Watermarked?
GPT-4o-mini -0.005 0.018 -5.984 No
GPT-4o 0.017 0.020 -4.211 No
GPT-3.5-turbo 0.028 0.030 -2.362 No
Gemini-1.5-flash 0.027 0.049 -1.474 No
Gemini-1.5-pro 0.018 0.038 -2.135 No
Our experimental results provide strong evidence that current closed-source model APIs do not contain watermarks. However, it is important to note that a key limitation of this experiment is our inability to verify ground truth labels, making it impossible to definitively confirm the accuracy of our detection results.
Appendix M Reversion Key Calculation of Water-Bag
In this section, we provide detailed calculations for determining reversion keys that satisfy the constraints in Equation 11. Let $p=P_{M}(y_{i}|x,y_{1:i-1})$ represent the original model distribution, and $q=F(p,f(K_{j},y_{i-n:i-1}))$ represent the distribution after modification using key $K_{j}$.
According to Equation 11, we have:
$\frac{1}{2}(q+F(p,f(\overline{K_{j}},y_{i-n:i-1})))=p$ (31)
Through algebraic manipulation, we can derive the required modification for the reversion key:
$F(p,f(\overline{K_{j}},y_{i-n:i-1}))=2p-q$ (32)
This equation provides the concrete method for calculating the reversion key $\overline{K_{j}}$. Specifically, for any input sequence $y_{i-n:i-1}$, the function $f(\overline{K_{j}},y_{i-n:i-1})$ must map the original distribution $p$ to $2p-q$ to satisfy Equation 11.
It is important to note that a reversion key need not be restricted to numerical values. Any key that produces the required distributional modification qualifies as a valid reversion key, as long as it accurately satisfies the constraint equation.